# Detection of disk-jet co-precession in a tidal disruption event

### v1

@@ 1.2.4 Extended ROentgen Survey with an Imaging Telescope Array (eROSITA)/Spectrum-Roentgen-Gamma (SRG)
We used eROSITA (?) all-sky survey (eRASS1) data, where our source is located in the Skytile 050087https://erosita.mpe.mpg.de/dr1/erodat/skyview/skytile/50087. This region was observed between February 3 and 16, 2020, nearly four years before the TDE. We analyzed the event file using the eROSITA Science Analysis Software System (eSASS) following the official recommended cookbookhttps://erosita.mpe.mpg.de/dr1/eSASS4DR1/eSASS4DR1_cookbook. Our source can be significantly identified, and we extracted its spectrum using the srctool task. The spectrum was re-binned using the grppha task, with mincounts=3.
@@ 1.3.1 Very Large Array (VLA)
The radio data were reduced using the Common Astronomy Software Application package (CASA 6.5.4) (?, ?), following standard data reduction procedures. The calibrated measurement sets were downloaded from the NRAO archive§§§https://data.nrao.edu/portal/, which were calibrated using the automated VLA calibration pipelinehttps://science.nrao.edu/facilities/vla/data-processing/pipeline available in the CASA. Flux density and bandpass calibration were conducted using 3C48, 3C147, 3C286, or 3C138, while J0339–0146 was used as a gain calibrator. The observation taken on MJD 60395.8, which failed to pass the pipeline’s quality assessment, was excluded from the analysis. Images of the AT2020afhd field were produced using the CASA task tclean with a Briggs robust parameter of 0.5 (?). We measured the flux density of AT2020afhd by fitting a point-source Gaussian model in the image plane with the CASA task imfit and determined the off-source rms using the CASA task imstat. The flux density error was estimated as the quadratic sum of the off-source rms, the fitting error from imfit, and a typical systematic error of 5% of the source flux density. To construct the lightcurve, all C-band observations, including 2 multi-band epochs, were imaged using a reference frequency at 5.5 GHz. The in-band spectral index at C band (VLA-$\alpha$) was calculated by imaging each spectral window (spw) separately for each epoch and fitting a power-law model with Python’s curve_fit package. To capture the spectral evolution, the broadband spectrum at the two multi-band epochs (i.e., observed on MJD 60437.9 and 60510.5) and three C-band epochs (i.e., observed on MJD 60399.8, 60461.7 and 60485.6) were constructed (see the selection criteria in the section ‘Radio long-term evolution’). Each set of 8 spws was imaged separately for the two multi-band epochs, and each set of 4 spws was imaged separately for three C-band epochs (see Fig. S5).
@@ 1.3.2 Australia Telescope Compact Array (ATCA)
ATCA conducted 32 observations of AT2020afhd from May 2 to November 12, 2024, as part of the DDT projects CX567, C3615, and CX591 (PI: Y. Wang). All the observations were observed at C band (4–10 GHz).
The radio data were reduced in CASA of the same version for the VLA data, following the tutorial https://casaguides.nrao.edu/index.php/ATCA_Tutorials and additionally using the ‘rflag’ and ‘tfcrop’ modes in the CASA task flagdata to flag the radio-frequency interference. In all observations, 1934–638 was used to calibrate the bandpass response, flux-density scale and polarization leakage, while 0336–019 was used to calibrate the time-variable complex gains. We imaged all the observations for the AT2020afhd field at C band with reference frequencies at 5.5 GHz and 9 GHz, each covering a bandwidth of 2048 MHz, using the CASA task tclean with natural weighting. The flux density and corresponding error measurements of AT2020afhd were the same as those of VLA. We calculated the spectral index for ATCA (i.e., ATCA-$\alpha$) using a two point (at 5.5 GHz and 9 GHz) power-law fitting.
@@ 1.3.3 Enhanced Multi-Element Radio Linked Interferometer Network (e-MERLIN)
We noted that four observations—RR17006_C_004_20240716, RR17006_C_006_20240723, RR18001_C_005_20240827 and RR18002_C_001_20241105—suffered from excessive visibility weights, resulting in suboptimal images produced by the pipeline. To address this, we flagged these problematic visibilities and re-imaged these observations using the CASA task tclean with the same parameters as the pipeline. RR18002_C_004_20241117 was excluded in the analysis, as no significant signals larger than 3 times off-source rms were detected at the target position in the final image. This was likely due to the majority ($\sim 85\%$) of the data being flagged and the short exposure ($\sim 2.5$ hours on target). The flux density and corresponding error measurements of AT2020afhd followed the same methodology as those for the VLA, including a 10% systematic error as suggested by the e-MERLIN team.
@@ 1.3.4 Very Long Based Array (VLBA)
We conducted four VLBA observations of AT2020afhd: on 2024-06-21, 2024-09-03, and 2024-11-01 at C band, and on 2024-06-29 at U band, with at least 8 participating antennas (10 for the first epoch). All observations were part of the DDT project BW161 (PI: Y. Wang). The VLBA data were correlated at the Array Operations Center in Socorro, USA. We phase-referenced our observations to the calibrator J0315–0151 with cycle time of 5 min (1 min on calibrator and 4 on target) for a total of $\sim$1.5 h each epoch for C band (6.2 GHz) and U band (15 GHz). We used the source J0339-0146 as fringe finder and bandpass calibrator. We performed standard a-priori gain calibration using the measured gains and system temperatures of each antenna. The data inspection and flagging, full calibration were done within the National Radio Astronomy Observatory Astronomical Image Processing System (?). We also corrected for ionosphere effects and source-structure effects of the phase-reference source in each band (?). After the calibration, 20% of the data from the second and third epochs were removed due to poor quality, resulting in noisier datasets compared to the first and fourth epochs.
@@
The observation log for all the radio epochs at C-band are available on Zenodo. The columns in the log include the observation date, time, MJD, radio telescope, telescope configuration, reference frequency of image, bandwidth, major axis of beam, minor axis of beam, position angle (PA) of beam, AT2020afhd’s flux density, AT2020afhd’s flux density error, off-source rms, and SNR (flux density divided by off-source rms) respectively.
@@ 1.3.5 Evaluation of calibration differences among radio arrays
To evaluate the flux density offset between VLA and ATCA, we used the only phase calibrator shared by both arrays, J0339–0146 (0336–019). For each observation, we inspected the output CASA log of the fluxscale task to get the fitted spectrum model of J0339–0146 and measured its flux density at 5.5 GHz. The results showed that J0339–0146 remained relatively stable during the studied periods, with mean flux densities of $\mu=1.30$ Jy (VLA) and $1.59$ Jy (ATCA), and standard deviations of $\sigma=0.11$ Jy for both arrays. This yielded a small coefficient of variation (CV), $\sigma/\mu\sim 6.9-8.5\%$, which is significantly lower than those observed in AT2020afhd, where $\sigma/\mu\sim 37.5-41.7\%$ and $\sim 25.0-33.3\%$ were measured before and after detrending the long-term evolution using Equation S2, respectively. The offset between the mean flux densities of the two arrays is approximately $18-22\%$.
Similarly, we selected the bandpass calibrator of our e-MERLIN observations, namely J1407+2827 (OQ208), to evaluate the calibration offsets between the VLA and e-MERLIN arrays. During the studied period, we identified five VLA observations of J1407+2827 at C band using B or BnA configurations from the NRAO archive. These observations utilized the same flux calibrator (3C286) with our e-MERLIN observations, and also demonstrated a partial overlap in UV coverage with our e-MERLIN observations. The five VLA epochs were taken on MJD 60455.2, 60466.0, 60484.3, 60554.680, and 60589.6. These observations were calibrated with CASA 6.5.4 following standard data reduction procedures. We found that J1407+2827 showed comparable amplitude within the same frequency range across the overlapping UV coverage for both arrays. Additionally, we derived the flux density of J1407+2827 at 5.07 GHz for each observation using the same method described earlier. Overall, the mean flux densities were $\mu=1.60$ Jy and $1.49$ Jy, with standard deviations of $\sigma=0.01$ Jy and $0.12$ for VLA and e-MERLIN, respectively. This again yielded a small CV, $\sigma/\mu\sim 0.6-8.1\%$. Thus, despite the larger scatter in flux densities from the e-MERLIN observations, the flux calibrations between the two arrays were found to be marginally consistent.
In conclusion, the low CV values of J0339-0146 and J1407+2827 at the observed frequency confirmed that the radio variations of AT2020afhd were not caused by flux calibrators. The offsets in the mean flux densities of the two calibrators, as determined from the three arrays, were up to 22%. To fully account for these calibration differences, we adopted the maximum offset value and reduced the ATCA flux density by 22% in Fig. 1 and throughout the analysis of the radio-X-ray cross-correlation function.
@@ 2 X-ray spectral analysis
X-ray spectral analysis was performed using XSPEC version 12.13.1 (?), with C-stat applied for goodness-of-fit testing. For XRT and NICER, the X-ray spectrum is ultra-soft, predominantly background-dominated above 2 keV, and is well described by an absorbed multi-color blackbody component; while for the XMM-Newton spectrum, where individual exposures are significantly longer than those of the other two facilities, we extended the fitting to the 0.3–10 keV energy band to assess any hard excess.
In addition to using tbabs_Gal to account for Galactic absorption, we introduced another component, tbabs_Loc, to evaluate the presence of local absorbers from the host galaxy. The value of tbabs_Gal was fixed at $4.72\times 10^{20}$ cm-2, as determined from the HI4PI survey (?). However, due to the narrow energy band and relatively low flux, tbabs_Loc could not be well constrained when fitting each spectrum individually. To robustly determine tbabs_Loc, we adopted a two-step approach: 1) jointly fitting spectra from the same facility with the data obtained before MJD 60530 when the target was in a high-luminosity phase, and comparing the results to ensure consistency; 2) jointly fitting spectra obtained at high and low luminosities, respectively, to evaluate whether absorption contributes to the observed modulation. The initial model applied was tbabs_Gal*zashift*tbabs_Lcol*diskbb for the XRT and NICER data, and tbabs_Gal*zashift*tbabs_Lcol*thcomp*diskbb for the XMM-Newton data. The redshift was fixed at 0.027. The abundance table and the photoelectric cross-section for tbabs were set to wilm (?), and vern (?), respectively. The unabsorbed flux of the diskbb component was calculated using the cflux command in the 0.3–2 keV.
@@
An absorption feature around 0.9 keV was observed in some of the observations from the three facilities spanning from July 13 to 27, suggesting the presence of clumpy outflows along the line of sight. A detailed analysis of these absorption lines will be presented in a forthcoming paper (Lin et al., in preparation).
@@
Unlike the above thermal blackbody spectrum, the SRG spectrum obtained prior to the TDE exhibits a powerlaw profile with a photon index of $2.07\pm 0.57$. The corresponding 0.3–2 keV luminosity is $10^{41.8\pm 0.1}\rm~erg~s^{-1}$, which is higher than the dip phase observed after 300 days since the re-brightening, when the X-ray emission experienced the second rapid drop. This suggests that the weak AGN activity observed in 2020 should have quenched by the time the TDE occurred.
3 Lomb-Scargle periodogram and cross-correlation function
The X-ray lightcurve was derived from the unabsorbed flux of the diskbb component in the 0.3–2 keV band, revealing clear quasi-periodic variations between August 3 and October 21, 2024, spanning a total of 79 days. To analyze these variations, we computed the Lomb-Scargle periodogram (LSP) (?) using the astropy library (?) with data from this 79-day period. To more accurately determine the period of the variations, we divided the NICER observations into individual GTIs with durations exceeding 300 seconds. The flux for each NICER GTI was measured following the method described in Section 1. Our analysis was performed on a GTI basis for the NICER data and on an observation basis for the Swift and XMM-Newton data. Given the average sampling interval of 1 day and the total duration of 79 days, we conducted the LSP analysis over a period search range of 1 to 100 days, similar to the range used in the study of AT2020ocn, which had a period of 15 days (?). This analysis identified a period of 19.6 days with a full width at half maximum (FWHM) of 3.0 days, as illustrated in the Fig. 4a.
To assess the statistical significance of the detected period, we first quantified the contribution of the LSP continuum by modeling it as a power-law, $P(f)\propto f^{\alpha}$. The fit was performed on the observed LSP, excluding the period range corresponding to the detected period signal, as defined by its FWHM. The best-fitting power-law index was determined to be $\alpha=-0.02\pm 0.05$. This value is consistent with 0 within the $1\sigma$ uncertainty range, indicating that the continuum was not significantly affected by red noise but was instead dominated by white noise.
To further test whether the continuum was consistent with white noise, we applied the algorithms described in (?). We compared the cumulative distribution function (CDF) and probability density function (PDF) of the LSP power values to those expected for white noise. The expected CDF followed $1-\exp(-z)$ (?), where $z$ represented LSP powers. The comparisons of the observed and expected CDF and PDF are shown in Figs. S4a and b. To quantify these comparisons, we conducted Kolmogorov-Smirnov (K-S) and Anderson-Darling (A-D) goodness-of-fit tests on the observed and expected CDF using the scipy library (?). The tests provided a K-S statistic of 0.0391 (p-value = 0.93) and an A-D statistic of 0.4278, which was well below the 10% critical value of 1.07. These results indicate that, at a 90% confidence level, the null hypothesis that the LSP power followed the expected white noise distribution could not be rejected Additionally, we performed 100,000 Monte Carlo simulations of the LSP continuum under the assumption of white noise and computed their corresponding K-S and A-D statistics. The distributions of these statistics are presented in EDFs. S4c and d. The K-S and A-D statistic values of the observed CDF fell within the $1\sigma$ range of the distribution of the simulated white noise statistic values. These findings robustly indicated that the LSP continuum was statistically consistent with white noise.
To investigate the correlation between radio and X-ray variations of AT2020afhd, we used the Python-based discrete cross-correlation function (PyCCF, (?)), focusing on data collected between June 19 and October 21, 2024. The analysis revealed that the cross-correlation function (CCF) peaked at a time lag of $\sim-19.0$ days (see Fig. 4b). To assess the global significance of this correlation, we generated 100,000 random radio lightcurves, preserving the same observation sampling while randomizing the luminosities within the observed minimum and maximum values. We tested the CCF across time lags spanning 45-45 to $10$ days, a range covering $\pm 1$ cycle of the periodic signal. Among these simulations, only 2 spurious CCF signals were detected, corresponding to a statistical significance of approximately $4.26\sigma$. Furthermore, we estimated the uncertainties in the time lag by focusing on the range of 30-30 to 10-10 days, corresponding to a single period cycle around the CCF peak. Using 100,000 Monte Carlo simulations with a bin size of 0.1 days and bootstrapping methods to estimate uncertainties (?, ?), we determined a time lag of $19.0^{+0.7}_{-0.6}$ days for the radio relative to the X-ray. This result is consistent with the observed period, indicating that the X-ray and radio emissions were synchronized. As described in Section 3.5, we applied a 22% offset to the observed ATCA flux density. Using the original ATCA flux density and following the same methodology outlined above, we obtained a statistical significance of $3.85\sigma$ and a time lag of $19.1^{+0.6}_{-0.7}$ days. These results indicate that the calibration offset had no significant impact on the derived time lag or the statistical significance of the correlation.

### v2

@@ 1.2.4 Extended ROentgen Survey with an Imaging Telescope Array (eROSITA)/Spectrum-Roentgen-Gamma (SRG)
We used eROSITA (?) all-sky survey (eRASS1) data, where our source is located in the Skytile 050093 (https://erosita.mpe.mpg.de/dr1/erodat/skyview/skytile/050093/?goto=48.39944,-2.15094). This region was observed between February 3 and 16, 2020, nearly four years before the TDE. We analyzed the event file using the eROSITA Science Analysis Software System (eSASS) following the official recommended cookbook
(https://erosita.mpe.mpg.de/dr1/eSASS4DR1/eSASS4DR1_cookbook/). Our source can be clearly identified, and we extracted its spectrum using the srctool task. The spectrum was re-binned using the grppha task, with mincounts=3.
@@ 1.3.1 Very Large Array (VLA)
The radio data were reduced using the Common Astronomy Software Application package (CASA 6.5.4, (?, ?)), following standard data reduction procedures. The calibrated measurement sets were downloaded from the NRAO archive (https://data.nrao.edu/portal/), which were calibrated using the automated VLA calibration pipeline (https://science.nrao.edu/facilities/vla/data-processing/pipeline) available in the CASA. Flux density and bandpass calibration were conducted using 3C48, 3C147, 3C286, or 3C138, while J0339–0146 was used as a gain calibrator. The observation taken on MJD 60395.8, which failed to pass the pipeline’s quality assessment, was excluded from the analysis. Images of the AT2020afhd field were produced using the CASA task tclean with a Briggs robust parameter of 0.5 (?). We measured the flux density of AT2020afhd by fitting a point-source Gaussian model in the image plane with the CASA task imfit and determined the off-source rms using the CASA task imstat. The flux density error was estimated as the quadratic sum of the off-source rms, the fitting error from imfit, and a typical systematic error of 5% of the source flux density. To construct the lightcurve, all C-band observations, including 2 multi-band epochs, were imaged using a reference frequency at 5.5 GHz. The in-band spectral index at C band (VLA-$\alpha$) was calculated by imaging each spectral window (spw) separately for each epoch and fitting a power-law model with Python’s curve_fit package. To capture the spectral evolution, the broadband spectrum at the two multi-band epochs (i.e., observed on MJD 60437.9 and 60510.5) and three C-band epochs (i.e., observed on MJD 60399.8, 60461.7 and 60485.6) were constructed (see the selection criteria in Supplementary 3). Each set of 8 spws was imaged separately for the two multi-band epochs, and each set of 4 spws was imaged separately for the three selected C-band epochs (see fig. S4).
@@ 1.3.2 Australia Telescope Compact Array (ATCA)
The ATCA conducted 32 observations of AT2020afhd between May 2 and November 12, 2024, as part of the DDT projects CX567 and CX591, and the Non A-Priori Assignable project C3615 (PI: Y. Wang). All the observations were observed at C band (4–10 GHz).
The radio data were reduced in CASA of the same version for the VLA data, following the tutorial (https://casaguides.nrao.edu/index.php/ATCA_Tutorials) and additionally using the ‘rflag’ and ‘tfcrop’ modes in the CASA task flagdata to flag the radio-frequency interference. In all observations, 1934–638 was used to calibrate the bandpass response, flux-density scale and polarization leakage, while 0336–019 was used to calibrate the time-variable complex gains. We imaged all the observations for the AT2020afhd field at C band with reference frequencies at 5.5 GHz and 9 GHz, each covering a bandwidth of 2048 MHz, using the CASA task tclean with natural weighting. The flux density and corresponding error measurements of AT2020afhd were the same as those of VLA. We calculated the spectral index for ATCA (i.e., ATCA-$\alpha$) by fitting a power-law to the two-point flux measurements at 5.5 GHz and 9 GHz.
@@ 1.3.3 Enhanced Multi-Element Radio Linked Interferometer Network (e-MERLIN)
We noted that four observations—RR17006_C_004_20240716, RR17006_C_006_20240723, RR18001_C_005_20240827 and RR18002_C_001_20241105—suffered from excessive visibility weights, resulting in suboptimal images produced by the pipeline. To address this, we flagged these problematic visibilities and re-imaged these observations using the CASA task tclean with the same parameters as the pipeline. RR18002_C_004_20241117 was excluded in the analysis, as no significant signal, i.e., greater than three times the off-source rms, was detected at the target position in the final image. This was likely due to the majority ($\sim$85%) of the data being flagged and the short exposure ($\sim$2.5 hours on target). The flux density and corresponding error measurements of AT2020afhd followed the same methodology as those for the VLA, including a 10% systematic error as suggested by the e-MERLIN team.
@@ 1.3.4 Very Long Based Array (VLBA)
We conducted four VLBA observations of AT2020afhd: on 2024-06-21, 2024-09-03, and 2024-11-01 at C band, and on 2024-06-29 at U band, with at least 8 participating antennas (10 for the first epoch). All observations were part of the DDT project BW161 (PI: Y. Wang). The VLBA data were correlated at the Array Operations Center in Socorro, USA. We phase-referenced our observations to the calibrator J0315–0151 with cycle time of 5 min (1 min on calibrator and 4 min on target) for a total of $\sim$1.5 h each epoch for C band (6.2 GHz) and U band (15 GHz). We used the source J0339-0146 as fringe finder and bandpass calibrator. We performed standard a-priori gain calibration using the measured gains and system temperatures of each antenna. The data inspection and flagging, full calibration were done within the National Radio Astronomy Observatory Astronomical Image Processing System (?). We also corrected for ionosphere effects and source-structure effects of the phase-reference source in each band (?). After the calibration, 20% of the data from the second and third epochs were discarded due to poor quality, resulting in noisier datasets compared to the first and fourth epochs.
@@
The observation log for all the radio epochs at C-band are available on Zenodo. The columns in the log include the observation date, time, MJD, radio telescope, telescope configuration, reference frequency of image, bandwidth, major axis of beam, minor axis of beam, position angle (PA) of beam, flux density, flux density error, off-source rms, and SNR (flux density divided by off-source rms), respectively.
@@ 1.3.5 Evaluation of calibration differences among radio arrays
To evaluate the flux density offset between VLA and ATCA, we used the only phase calibrator shared by both arrays, J0339–0146 (0336–019). For each observation, we inspected the output CASA log of the fluxscale task to get the fitted spectrum model of J0339–0146 and measured its flux density at 5.5 GHz. The results showed that J0339–0146 remained relatively stable during the studied periods, with mean flux densities of $\mu=1.30$ Jy (VLA) and $1.59$ Jy (ATCA), and standard deviations of $\sigma=0.11$ Jy for both arrays. This yielded a small coefficient of variation (CV), $\sigma/\mu\sim 6.9$–8.5%, which is substantially lower than those observed in AT2020afhd, where $\sigma/\mu\sim 37.5$–41.7% and $\sim 25.0$–33.3% were measured before and after detrending the long-term evolution using equation S2, respectively. The offset between the mean flux densities of the two arrays is approximately 18–22%.
Similarly, we selected the bandpass calibrator of our e-MERLIN observations, namely J1407+2827 (OQ208), to evaluate the calibration offsets between the VLA and e-MERLIN arrays. During the studied period, we identified five VLA observations of J1407+2827 at C band using B or BnA configurations from the NRAO archive. These observations utilized the same flux calibrator (3C286) with our e-MERLIN observations, and also demonstrated a partial overlap in UV coverage with our e-MERLIN observations. The five VLA epochs were taken on MJD 60455.2, 60466.0, 60484.3, 60554.680, and 60589.6. These observations were calibrated with CASA 6.5.4 following standard data reduction procedures. We found that J1407+2827 showed comparable amplitude within the same frequency range across the overlapping UV coverage for both arrays. Additionally, we derived the flux density of J1407+2827 at 5.07 GHz for each observation using the same method described earlier. Overall, the mean flux densities were $\mu=1.60$ Jy and $1.49$ Jy, with standard deviations of $\sigma=0.01$ Jy and $0.12$ for VLA and e-MERLIN, respectively. This again yielded a small CV, $\sigma/\mu\sim 0.6$–8.1%. Thus, despite the larger scatter in flux densities from the e-MERLIN observations, the flux calibrations between the two arrays were found to be marginally consistent.
In conclusion, the low CV values of J0339-0146 and J1407+2827 at the observed frequency confirmed that the radio variations of AT2020afhd were not caused by flux calibrators. The offsets in the mean flux densities of the two calibrators, as determined from the three arrays, were up to 22%. To fully account for these calibration differences, we applied the maximum offset value of 22% to reduce the ATCA flux densities, and used this correction consistently throughout the analysis of the radio-X-ray cross-correlation function.
@@ 2 X-ray spectral analysis
X-ray spectral analysis was performed using XSPEC version 12.13.1 (?), with C-stat applied for goodness-of-fit testing. For XRT and NICER, the X-ray spectrum is ultra-soft, predominantly background-dominated above 2 keV, and is well described by an absorbed multi-color blackbody component; while for the PN spectrum, where individual exposures are significantly longer than those of the other two facilities, we extended the fitting to the 0.3–10 keV energy band to assess any hard excess.
In addition to using tbabs_Gal to account for Galactic absorption, we introduced another component, tbabs_Loc, to evaluate the presence of local absorbers from the host galaxy. The value of tbabs_Gal was fixed at $4.72\times 10^{20}$ cm-2, as determined from the HI4PI survey (?). However, due to the narrow energy band and relatively low flux, tbabs_Loc could not be well constrained when fitting each spectrum individually. To robustly determine tbabs_Loc, we adopted a two-step approach: 1) jointly fitting spectra from the same facility with the data obtained before MJD 60530 when the target was in a high-luminosity phase, and comparing the results to ensure consistency; 2) jointly fitting spectra obtained at high and low luminosities, respectively, to evaluate whether absorption contributes to the observed modulation. The initial model applied was tbabs_Gal*zashift*tbabs_Lcol*diskbb for the XRT and NICER data, and tbabs_Gal*zashift*tbabs_Lcol*thcomp*diskbb for the PN data. The redshift was fixed at 0.027. The abundance table and the photoelectric cross-section for tbabs were set to wilm (?), and vern (?), respectively. The unabsorbed flux of the diskbb component was calculated using the cflux command in the 0.3–2 keV.
@@
An absorption feature around 0.9 keV was observed in some of the observations from the three facilities spanning from July 13 to 27, suggesting the presence of clumpy outflows along the line of sight. A detailed analysis of these absorption lines is presented in a separate study (?).
@@
Unlike the above thermal blackbody spectrum, the SRG spectrum obtained prior to the TDE exhibits a powerlaw profile with a photon index of $2.07\pm 0.57$. The corresponding 0.3–2 keV luminosity is $10^{41.8\pm 0.1}\rm~erg~s^{-1}$, which exceeds the luminosity measured after day 300 following the re-brightening, when the X-ray emission underwent a second rapid decline. This suggests that the AGN activity observed in 2020 had likely weakened by the time the TDE occurred.